Let f(x) = -x - 5 and g(x) = 2x + 3 Find the indicated value.
f(x) when x = -10
Any help with the given equation would Greatly appreciated.

Answer by stanbon(75887)

Let f(x) = -x - 5
Find f(x) when x = -10
----------------
f(-10) = -(-10)-5
f(-10) = 10-5 = 5
====================
Chers,
Stan H.